OpenStudy (anonymous):

The solutions to a certain quadratic equation are 3 and −3. Find the equation.

OpenStudy (anonymous):

x^2=9

OpenStudy (anonymous):

It means that x=3 or x=-3 will give you 0: (x-3)(x+3)=0 x^2-9=0